On July 15, 2015, Tina declared and issued a 15 percent stock dividend. Prior to this dividend, Tina had 50,000 shares of $10 par value common stock issued and outstanding. The market value of Tina’s common stock on July 15, 2015, was $23 per share. All else equal, what would we expect the approximate market value per share of Tina’s stock to be after the dividend?
Answer #1

Number of shares = 50000 shares

Value of equity = 50000 * 23 = 1,150,000

stock dividend issued = 50000 * 15% = 7500 shares

Market value of shares = 1,150,000 / 57500 = $20 per share